Ticket Number: HEA-NOC/20080903-10 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080903 11:55 UTC+1 Problem Start: 20080903 12:30 UTC+1
Ticket Update: 20080903 14:03 UTC+1 Problem End:
Migration of ESRI's current connections to HEAnet network
No service interruption as this is a new client that is migrating to
HEAnet. A router is to be installed at the ESRI on Wednesday 3rd Sept at
12:30pm. When completed ESRI's traffic will be routed via the HEAnet
20080903 11:53 GL Confirmed with ESRI about install.
20080903 14:02 GL Transition completed. ESRI's traffic will be routed via
the HEAnet network. Will monitor for 24 hours
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080903-10
HEAnet Limited. Registered in Ireland, No. 275301.